60% of people those who purchase life insurance are women. If the owners of 5 life


insurance are randomly selected, then find the probability that there are


(i) no women


(ii) exactly 3 women


(iii)more than 2 women.


1
Expert's answer
2021-12-24T08:33:14-0500

Solution:

"n=5, p=60\%=0.6, q=0.4
\\X\sim Bin(n,p)"

(i)

"P(X=0)=\ ^5C_0(0.6)^0(0.4)^5=0.01024"

(ii)

"P(X=3)=\ ^5C_3(0.6)^3(0.4)^2=0.3456"

(iii)

"P(X>2)=1-P(X\le 2)
\\=1-[P(X=0)+P(X=1)+P(X=2)]
\\=\ ^5C_0(0.6)^0(0.4)^5+\ ^5C_1(0.6)^1(0.4)^4+\ ^5C_2(0.6)^2(0.4)^3
\\=0.31744"